# The asymptoticity of extremal length in Teichmüller space

### v1

@@ 5.2. The detour metric
Walsh [21] showed that for any $X\in\mathcal{T}(S)$ and $\xi\in\partial_{B}\mathcal{T}(S)$, there exists a unique Teichmüller ray starting from $X$ and converging to $\xi$. We denote by $\mathcal{B}_{q}$ the Busemann point in $\partial_{B}\mathcal{T}(S)$, which is the limit of a Teichmüller ray $\mathcal{R}_{q,X}(t)$. Walsh also established a formula of detour distance between two Busemann points which are the limits of two Teichmüller rays from the basepoint of $\overline{\mathcal{T}(S)}^{hor}$ (see the Corollary $3$ in [21]). Amano [2] generalized this formula to any two Busemann points associated with distinct Teichmüller rays originating from different satrting points.
Proposition 5.1 ([2]).

### v2

@@ 5.2. The detour metric
Walsh [21] showed that for any $X\in\mathcal{T}(S)$ and $\xi\in\partial_{B}\mathcal{T}(S)$, there exists a unique Teichmüller ray starting from $X$ and converging to $\xi$. We denote by $\mathcal{B}_{q}$ the Busemann point in $\partial_{B}\mathcal{T}(S)$, which is the limit of a Teichmüller ray $\mathcal{R}_{q,X}(t)$. Walsh also established a formula of detour distance between two Busemann points which are the limits of two Teichmüller rays from the basepoint of $\overline{\mathcal{T}(S)}^{hor}$ (see [21, Corollary 3]). Amano [2] generalized this formula to any two Busemann points associated with distinct Teichmüller rays originating from different satrting points.
Proposition 5.1 ([2, Proposition 4.14]).